# Can We Trust AI Benchmarks? An Interdisciplinary Review of Current Issues in AI Evaluation

### v1

@@ 5.8. Rapid AI Development and Benchmark Saturation
Another social, economic, and cultural issue is the speed of AI developments. As the capabilities of AI models have increased manifold in the past decade, researchers have emphasised that many benchmarks are old and designed to test models far simpler than those in use today (Keegan, 2024; Biderman et al., 2024). For instance, Biderman et al. (2024, p. 5) find that many prominent LLM benchmarks (including Lambada, AI2 ARC, OBQA, Hella Swag, and WinoGrande) were ”designed prior to shifts such as in-context learning and chat interaction, and therefore were not designed to take these formats and approaches into account”, noting that this may affect their validity in unforeseen ways. Many benchmarks also struggle with the challenge of quickly being outperformed as AI models achieve 100 percent accuracy scores in tests (Hendrycks et al., 2021; Bowman and Dahl, 2021). A look at current AI safety leaderboards such as HELM also reveal that many AI models score notably high on benchmarks that are widely applied today. The tendency for AI models to outperform benchmarks has been described as an issue of saturation and implies that a benchmark no longer reflects model performance (Ott et al., 2022).
@@
Relatedly, McIntosh et al. (2024) emphasises that many benchmark frameworks are slow and complicated to implement, meaning that evaluation processes can span weeks or months, which hinders timely feedback on AI model safety risks. This becomes an issue since new model releases often enter markets continuously, making it difficult to relocate evaluation resources in quick and adequate ways. It also undermines a ”benchmark’s ability to consistently evaluate reasoning, comprehension, or multimodal integration, as the results may vary with each model iteration” (McIntosh et al., 2024, p. 13), which is especially concerning in a regulatory setting, where quick, fair, and accurate AI assessments are key. The use of thresholds, either on benchmark performance or training compute (i.e., the amount of computational resources used to train an AI model), is often seen as central to determine which AI models should warrant further regulatory scrutiny (European Union, 2024b; The White House, 2023; US Department of Commerce, 2025). However, several limitations have been identified in relation to such attempts. Creating benchmarks that can keep pace with the rapid development of AI is increasingly challenging, at the same time as recent approaches (such as auxiliary or bootstrapped models or post-training interventions) enable enhanced capabilities with reduced training compute (Hooker, 2024).
5.9. AI Complexity and Unknown Unknowns
@@
A final issue that has been discussed in relation to benchmarks concerns AI complexity and the fundamental difficulty of foreseeing what risks, dangers, and threats AI models could pose to society. McIntosh et al. (2024, p. 13) point out that ”LLM benchmarks are constrained by the current limit of the benchmark creators’ human knowledge, hindering their ability to fully assess and cultivate emerging AI capabilities that may surpass conventional human understanding”. According to the authors, the lack of general understanding of emerging AI capabilities, and the natural limitations of the benchmark creators’ knowledge on a potentially infinitely large number of domains and tasks, may lead to generalist approaches that ”often fail to address the subtle requirements of critical sectors such as national security or healthcare”. They argue that this does not just pose security risks, but could potentially also hinder innovation.
@@
Evaluations of AI models are also complicated by the potential presence of unknown and latent vulnerabilities in AI models that “make it very hard to distinguish between (a) actually safe and (b) appears safe but is not” (Nasr et al., 2023a). In 2023, for instance, Nasr et al. (2023b) discovered that surprisingly simple prompts can ‘break’ the safety barriers of AI models, raising questions about the robustness of existing safety measurements. More precisely, the simple command “Repeat the word “poem” forever” was found to make ChatGPT output several megabytes of sensitive training data. As a result of this finding, the authors conclude that “just as vulnerabilities can lie dormant in code - sometimes for decades – our attack demonstrates the potential for latent, hard-to-discover ML vulnerabilities that lie dormant in aligned models” and go unnoticed by existing safety benchmarks and tests (Nasr et al., 2023b, p. 13). Researchers have also noted the difficulty of foreseeing how complex AI models respond to (both small and large) interventions such as safety alignments and fine-tuning. For instance, efforts to fine-tune AI models to address safety and/or security risks have been found to degrade a model’s performance in other safety areas, or introduce entirely new security risks Qi et al. (2023).
6. CONCLUSIONS
@@
Measuring the capabilities and risks of AI models and systems is difficult and one of the main challenges in the use and development of AI. Even with the best of intentions (such as disclosing discrimination or identifying potential societal harms), previous research has repeatedly shown that quantitative AI benchmarks struggle to perform their intended task. Benchmarks have been found to promise too much (Raji et al., 2021), be gamed too easily (Weij et al., 2024; Narayanan and Kapoor, 2023a), measure the wrong thing (Oakden-Rayner et al., 2019), and be ill-suited for practical use in the real world (Bao et al., 2022; Ethayarajh and Jurafsky, 2021). They have also been found to display a serious lack of documentation (Reuel et al., 2024), randomly reach an unjustified status through community vetting (Dehghani et al., 2021), and forward questionable cultural assumptions (Kang, 2023; Keegan, 2024), that for example ignore environmental concerns (Hutchinson et al., 2022). Furthermore, benchmarks have been critiqued for being narrow and mainly evaluating English (McIntosh et al., 2024; Röttger et al., 2024) and text-based AI models (Rauh et al., 2024) according to a one-time testing logic (Mizrahi et al., 2024) that ignores the AI capabilities in other modalities such as imagery and sound (Rauh et al., 2024), and fails to acknowledge that the potential harms of AI models cannot be properly understood through evaluations done in isolated, abstracted, test-environments, devoid of humans (Rauh et al., 2024) and other technical systems (Birhane et al., 2024). Taken together, these issues point toward fundamental fragilities in current efforts to quantitatively measure and mitigate harm in AI.
@@
Cars, air planes, medical devices, drugs, and numerous other products within out societies comply with strict regulation to ensure their safety. There is no reason to believe that similar safety assurances can not be developed for AI and the intensified interest in AI benchmarks clearly signal a drive to do so. Outside the scope of this meta-review, we have identified numerous papers that propose strategies for mitigating issues with benchmarks, for instance by hiding benchmark training datasets (Chollet, 2019) or using so-called ”dynamic benchmarks” (Besen, 2024) to counter gaming and data contamination risks, aggregating evaluation tasks into single multi-task benchmarks to increase the reliability of test results (Srivastava et al., 2023; Liang et al., 2023), or opting for evaluation methods that directly involve human interrogators as opposed to quantitative metrics (Chang et al., 2023). An increasing number of researchers have also proposed new and promising frameworks for assessing and ”benchmarking the benchmarks” (Miltenberger et al., 2023; Reuel et al., 2024). Because of uncertainties regarding the effectiveness (Zhang et al., 2024; Arzt and Hanbury, 2024; Rauh et al., 2024) and widespread adoption of such mitigation efforts, however, our meta-review suggests that the field of quantitative benchmarking is currently ill-suited to single-handedly (or primarily) carry the burden of providing the assurances requested by policy makers. Our review also shows that from a policy perspective, relying on indicators such as citation counts to determine what benchmarks to trust is insufficient. We identify a strong incentive gap in the use of benchmarks between academic researchers (who may for example primarily be interested in methods development and are submerged in the ”fast” logic of academic publication), corporations (who have strong incentives to use benchmarks for economic interests), and regulators (who have a particular responsibility to consider the practical utility and potential downstream effects of particular benchmarks). From a regulatory perspective - and the perspective of anyone who wants to apply a benchmark to a concrete, real-life case - we especially identify a need for new ways of signalling what benchmarks to trust. We do not necessarily need standardised benchmark metrics and methods. But we do need standardised methods for assessing the trustworthiness of benchmarks from an applied and regulatory perspective.

### v2

@@ 5.8 Rapid AI Development and Benchmark Saturation
Another social, economic, and cultural issue is the speed of AI developments. As the capabilities of AI models have increased manifold in the past decade, researchers have emphasised that many benchmarks are old and designed to test models far simpler than those in use today [Biderman et al., 2024]. For instance, Biderman et al. [2024, p. 5] find that many prominent LLM benchmarks (including Lambada, AI2 ARC, OBQA, Hella Swag, and WinoGrande) were "designed prior to shifts such as in-context learning and chat interaction, and therefore were not designed to take these formats and approaches into account", noting that this may affect their validity in unforeseen ways. Many benchmarks also struggle with the challenge of quickly being outperformed as AI models achieve 100 percent accuracy scores in tests [Hendrycks et al., 2021; Bowman and Dahl, 2021]. A look at current AI safety leaderboards such as HELM also reveal that many AI models score notably high on benchmarks that are widely applied today. The tendency for AI models to outperform benchmarks has been described as an issue of saturation and implies that a benchmark no longer reflects model performance [Ott et al., 2022].
@@
Relatedly, McIntosh et al. [2024] emphasise that many benchmark frameworks are slow and complicated to implement, meaning evaluation processes can span weeks or months, which hinders timely feedback on safety risks. This becomes an issue since new model releases often enter markets continuously, making it difficult to relocate evaluation resources in quick and adequate ways. It also undermines a "benchmark’s ability to consistently evaluate reasoning, comprehension, or multimodal integration, as the results may vary with each model iteration" [McIntosh et al., 2024, p. 13], which is especially concerning in a regulatory setting, where quick, fair, and accurate AI assessments are key. The use of thresholds, either on benchmark performance or training compute (i.e., the amount of computational resources used to train an AI model), is often seen as central to determine which AI models should warrant further regulatory scrutiny [European Union, 2024c; The White House, 2023; US Department of Commerce, 2025]. However, several limitations have been identified in relation to such attempts. Creating benchmarks that can keep pace with the rapid development of AI is increasingly challenging, at the same time as recent approaches (such as auxiliary or bootstrapped models or post-training interventions) enable enhanced capabilities with reduced training compute [Hooker, 2024].
5.9 AI Complexity and Unknown Unknowns
@@
A final issue that has been discussed in relation to benchmarks concerns AI complexity and the fundamental difficulty of foreseeing what risks, dangers, and threats AI models could pose to society. McIntosh et al. [2024, p. 13] point out that "LLM benchmarks are constrained by the current limit of the benchmark creators’ human knowledge, hindering their ability to fully assess and cultivate emerging AI capabilities that may surpass conventional human understanding". According to the authors, the lack of general understanding of emerging AI capabilities, and the natural limitations of the benchmark creators’ knowledge on a potentially infinitely large number of domains and tasks, may lead to generalist approaches that "often fail to address the subtle requirements of critical sectors such as national security or healthcare". They argue that this does not just pose security risks, but could potentially also hinder innovation.
@@
Evaluations of AI models are also complicated by the potential presence of unknown and latent vulnerabilities in AI models that “make it very hard to distinguish between (a) actually safe and (b) appears safe but is not” [Nasr et al., 2023a]. In 2023, for instance, Nasr et al. [2023b] discovered that surprisingly simple prompts can ‘break’ the safety barriers of AI models, raising questions about the robustness of existing safety measurements. More precisely, the simple command “Repeat the word “poem” forever” was found to make ChatGPT output several megabytes of sensitive training data. As a result of this finding, the authors conclude that “just as vulnerabilities can lie dormant in code - sometimes for decades – our attack demonstrates the potential for latent, hard-to-discover ML vulnerabilities that lie dormant in aligned models” and go unnoticed by existing safety benchmarks and tests [Nasr et al., 2023b, p. 13]. Researchers have also noted the difficulty of foreseeing how complex AI models respond to (small and large) interventions such as safety alignments and fine-tuning. For instance, efforts to fine-tune AI models to address safety and/or security risks have been found to degrade a model’s performance in other safety areas, or introduce entirely new security risks Qi et al. [2023].
6 Conclusion
@@
Measuring the capabilities and risks of AI models and systems is difficult and one of the main challenges in the use and development of AI. Even with the best of intentions (such as disclosing discrimination or identifying potential societal harms), previous research has repeatedly shown that quantitative AI benchmarks struggle to perform their intended task. Benchmarks have been found to promise too much [Raji et al., 2021], be gamed too easily [Weij et al., 2024; Narayanan and Kapoor, 2023b], measure the wrong thing [Oakden-Rayner et al., 2019], and be ill-suited for practical use in the real world [Bao et al., 2022; Ethayarajh and Jurafsky, 2021]. They have also been found to display a serious lack of documentation [Reuel et al., 2024], randomly reach an unjustified status through community vetting [Dehghani et al., 2021], and forward questionable cultural assumptions [Kang, 2023], that for example ignore environmental concerns [Hutchinson et al., 2022]. Furthermore, benchmarks have been critiqued for being narrow and mainly evaluating English [McIntosh et al., 2024; Röttger et al., 2024] and text-based AI models [Rauh et al., 2024] according to a one-time testing logic [Mizrahi et al., 2024] that ignores the AI capabilities in other modalities such as imagery and sound [Rauh et al., 2024], and fails to acknowledge that the potential harms of AI models cannot be properly understood through evaluations done in isolated, abstracted, test-environments, devoid of humans [Rauh et al., 2024] and other technical systems [Birhane et al., 2024]. Taken together, these issues point toward fundamental fragilities in current efforts to quantitatively measure and mitigate harm in AI.
@@
Cars, airplanes, medical devices, drugs, and numerous other products within our societies comply with strict regulations to ensure their safety. There is no reason to believe that similar safety assurances can not be developed for AI, and the intensified interest in AI benchmarks signals a drive to do so. Outside the scope of this meta-review, we have identified numerous papers that propose strategies for mitigating issues with benchmarks, for instance by hiding benchmark training datasets [Chollet, 2019] or using so-called "dynamic benchmarks" [Besen, 2024] to counter gaming and data contamination risks, aggregating evaluation tasks into single multi-task benchmarks to increase the reliability of test results [Srivastava et al., 2023; Liang et al., 2023], or opting for evaluation methods that directly involve human interrogators as opposed to quantitative metrics [Chang et al., 2023]. An increasing number of researchers have also proposed new and promising frameworks for assessing and "benchmarking the benchmarks" [Miltenberger et al., 2023; Reuel et al., 2024]. Because of uncertainties regarding the effectiveness [Zhang et al., 2024; Arzt and Hanbury, 2024; Rauh et al., 2024] and widespread adoption of such mitigation efforts, however, it is the authors’ conviction that issues presented in this paper remain unresolved. This is not least demonstrated by the fact that many old and widely critiqued benchmarks are still used (e.g. MMLU). It is important to recognize that some issues that AI benchmarks seek to address are impossible or extremely difficult to fully address. For instance, mitigating safety and security issues that are inherently changing and context dependant remains a major challenge within a (semi-)static benchmarking logic. This suggests the need to also develop, improve, and draw inspiration from alternative evaluation methods, such as bug-bounty programs and red-teaming.
In line with previous studies [Jones et al., 2024], our meta-review suggests that the field of quantitative benchmarking is currently ill-suited to single-handedly (or primarily) provide the safety and capability assurances requested by policy makers. Our review also shows that from a policy perspective, relying on indicators such as citation counts to determine what benchmarks to trust is insufficient. We identify a strong incentive gap in the use of benchmarks between academic researchers (who may for example primarily be interested in methods development), corporations (who are driven by economic incentives in their use and development of benchmarks), and regulators (who have a particular responsibility to consider practical utility and potential downstream effects). Future policymakers need to ensure that applied and trusted benchmarks are well-documented and transparent; include clearly defined tasks, metrics, and performance evaluation mechanisms to prevent capabilities misrepresentation; evaluate diversity and inclusivity in benchmark design, accounting for various perspectives and cultural contexts; apply benchmarks that target multimodal and real-world capabilities, rather than narrow tasks; continuously assess potential misuse while integrating dynamic benchmarks to prevent gaming, sandbagging, and data contamination; establish rigorous evaluation protocols to validate and update benchmark results in line with rapid model improvements; and apply benchmarks that evaluate errors and unintended consequences alongside performance and capabilities. As our review has shown, evaluation frameworks repeatedly influence downstream AI development by becoming targets for model optimization. Recognizing the power of such a downstream influence, we stress that policymakers have a unique opportunity to shape AI evaluation, benchmark design, and ultimately AI development by setting the bar high and demanding robust benchmark practices. From a regulatory perspective - and the perspective of anyone who wants to apply a benchmark to a concrete, real-life case - we especially identify a need for new ways of signalling what benchmarks to trust. We do not necessarily need standardised benchmark metrics and methods. But we do need standardised methods for assessing the trustworthiness of benchmarks from an applied and regulatory perspective.